OpenStudy (anonymous):

find limx→π+ csc(2x) is it positive or a negative limit?

OpenStudy (anonymous):

\[\lim_{x \rightarrow \pi} \csc(2x)\] written like this ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

now you know the relationship between csc and sin ?

OpenStudy (anonymous):

yeah reciprocals right?

OpenStudy (anonymous):

it's called inverse actually : csc(x)=1/sin(x)

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

now csc(2x)=1/sin(2x)

OpenStudy (anonymous):

now sin(2x)=sin(x+x) which is....

OpenStudy (anonymous):

sin(a+b)=...

OpenStudy (anonymous):

continue

OpenStudy (anonymous):

yeah I'm lost

OpenStudy (anonymous):

sin(a+b)=?

OpenStudy (anonymous):

you have a formula for this

OpenStudy (anonymous):

sin(a+b)=sin(a)cos(b)+sin(b)cos(a)

OpenStudy (anonymous):

now sin(x+x)=...

OpenStudy (anonymous):

sin(2x)

OpenStudy (anonymous):

use the formula mate

OpenStudy (anonymous):

replace x with a and x with b and add that sum

OpenStudy (anonymous):

sin(a+b) ??

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

sin(2x)=2sinxcosx

OpenStudy (anonymous):

now

OpenStudy (anonymous):

your limit can be written in this form \[\lim_{x \rightarrow \pi}\frac{ 1 }{ 2\sin x \cos x }\]

OpenStudy (anonymous):

now replace x with pi and do the math...

OpenStudy (anonymous):

.027 ?

OpenStudy (anonymous):

thank you i got it